OpenStudy (anonymous):

What is the common denominator of 1/a + 1/b in the complex fraction (1/a-1/b ) / (1/a+1/b) a^2 b^2 a – b a + b ab

OpenStudy (freckles):

so it asked to find a common denominator, not the least common denominator?

OpenStudy (freckles):

can i ask you how you would combine the following fractions: \[\frac{1}{2}+\frac{1}{3}\]

OpenStudy (anonymous):

wouldnt you just add them straight across

OpenStudy (anonymous):

oh shoot or is the where you cant add the denominators unless theyre the same

OpenStudy (freckles):

you have to have find a common denominator to add the fractions one such being 6 2 goes into 6, 3 times 3 goes into 6, 2 times you can multiply the fraction fraction by 3/3 and the second fraction by 2/2 notice a common denominator can be found just by multiplying 2 and 3. \[\frac{1}{2}\frac{3}{3}+\frac{1}{3} \frac{2}{2} \\ \frac{3}{6}+\frac{2}{6} \text{ see we have found a common denominator } \\ \text{ now we can add } \\ \frac{3+2}{6}=\frac{5}{6} \\ \text{ so } \frac{1}{2}+\frac{1}{3}=\frac{5}{6}\]

OpenStudy (freckles):

so what do you think the a common denominator of 1/a,1/b would be so that you could add 1/a+1/b?

OpenStudy (freckles):

remember I just found a common denominator above just by multiplying the bottoms

OpenStudy (freckles):

what is the denominator of 1/a? what is the denominator of 1/b?

OpenStudy (anonymous):

ab?

OpenStudy (freckles):

yes a common denominator could be ab a common denominator can also be any multiple of that if you that ab*ab which is a^2b^2 is also a common denominator since this is multiple choice and you have both of those options i would go with the least common denominator ab